OpenStudy (anonymous):

y = 2x2 + 2 Inverse function

OpenStudy (akashdeepdeb):

Write x as y and y as x And then get the value of y in terms of x. Understood? :)

OpenStudy (akashdeepdeb):

y = 2x^2 + 2 x = 2y^2 + 2 \[y^2 = \frac{x-2}{2}\] \[y = \sqrt{\frac{x-2}{2}}\] Do all inverse questions like this. :)